OpenStudy (broskishelleh):

function f(x) = -16x2 + 128x shows the height of the rocket f(x), in feet, from the ground at time x seconds. What is the axis of symmetry of the graph of f(x), and what does it represent? x = 8; it takes 8 seconds to reach the maximum height and 8 seconds to fall back to the ground x = 4; it takes 4 seconds to reach the maximum height and 4 seconds to fall back to the ground x = 8; it takes 8 seconds to reach the maximum height and 16 seconds to fall back to the ground x = 4, it takes 4 seconds to reach the maximum height and 8 seconds

OpenStudy (broskishelleh):

I think A

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

any answer changes?

OpenStudy (broskishelleh):

No idea what you jst did? You added x, I don't see it

OpenStudy (anonymous):

Let me delete and redo it for you

OpenStudy (anonymous):

y= ax^2 + bx + c do you know this formula?

OpenStudy (broskishelleh):

Yep

OpenStudy (anonymous):

Good, next step this is the formula for finding axis of symmetry: axis of symmetry equation is x= -b/ 2a any questions?

OpenStudy (broskishelleh):

Nope

OpenStudy (anonymous):

Ok, now , look back at the equation f(x) = -16x2 + 128x a= -16 b= 128 plug a and b in to find it. So far so good.

OpenStudy (anonymous):

so plug a and b to the axis of symmetry equation x= -b/ 2a

OpenStudy (broskishelleh):

y= ax^2 + bx + c y= (-16)x^2 + (128)x + c

OpenStudy (anonymous):

Yes, that's right.

OpenStudy (broskishelleh):

x= -b/ 2a x= -(128)/ 2(-16)

OpenStudy (anonymous):

yes, correct.

OpenStudy (broskishelleh):

Solve it now?

OpenStudy (anonymous):

go ahead :)

OpenStudy (broskishelleh):

x= -b/ 2a x= -(128)/ 2(-16) x = -128/-32 x = 4

OpenStudy (anonymous):

perfect, so your answer changes to ?

OpenStudy (anonymous):

A, B ,C or D?

OpenStudy (broskishelleh):

B Or D

OpenStudy (anonymous):

look at this graph here http://www.skwirk.com/content/upload/images/Secondary/NSW/Year_10/Maths/graphing/tp3/ch2/tp3ch2_image1.jpg Notice that axis of symmetry divides the parabola into halves. Left side = right side.

OpenStudy (anonymous):

so, it takes 4 seconds to go up, 4 seconds to go down

OpenStudy (broskishelleh):

Yes I see what you mean

OpenStudy (anonymous):

so B, yes? no? :)

OpenStudy (broskishelleh):

So b yes